ECON 2301 CA2 Spring 2019 FPC Be sure to look at the instructors lecture notes and video over the PPF model if you are unclear about the questions. Marshalls Model Plane and Boat Productilon Planes go 80 70 60 50 40 30 20 10 Marshall Planes Boats 15 10 15 Boats 1. Given the above graph, calculate the following. (Be sure to show your work.) What is Marshalls opportunity cost of producing 10 boats? a. b. What is Marshalls opportunity cost of producing 30 planes? What is Marshalls marginal opportunity cost of producing boats? (In other words, for every one boat produced, how much does Marshall give up in plane production?) c. d. What is Marshalls marginal opportunity cost of producing planes?
Answer #1

a) opportunity cost of producing 10 boats= (90-30)/(10-0)= 6 planes

b) opportunity cost of producing 30 planes= (15-10)/(30-0)= 0.17

c) opportunity cost of producing boats= 90/15= 6 planes

d) opportunity cost of producing planes= 15/90= 0.17 boats.